The complement of an angle is twice the measure of the angle. What is the measure of each angle? Explain.
Answer by nyc_function(2741)

Let x = the angle
Let 90 - x = the complement of the angle

90 - x = 2x

90 = 2x + x

90 = 3x

90/3 = x

30 = x

The angles are 30 degrees and 60 degrees.